Phoebehall02:

To solve the equation 3sin(2x)=5cosx you should rewrite it as

AZ:

do you know the identity for sin(2x)

AZ:

\(\sin(2x) = 2\sin(x)\cos(x)\)

AZ:

How would you solve it now? You have a cos(x) on both sides. What can you simplify your equation to?